# Opinion Dynamics with Multiple Adversaries

### v1

@@ 2. Strategic Opinion Formation
Figure 3. Strategic misreports for the Polblogs dataset where $S$ is taken to be the top-50% of the agents in terms of their eigenvector centralities. The nodes are labeled either as liberal ($\bm{s}_{i}=-1$) or conservative ($\bm{s}_{i}=+1$), and we consider the nodes that change their beliefs as the nodes for which $\bm{z}_{i}^{\prime}$ and $\bm{z}_{i}$ do not have the same sign. In the scatterplots (a), (c), (d), (e), the shape of each point indicates whether that user changed belief or not, and the color indicates their true (intrinsic) opinion. Overall, we discover a higher amount of change among liberal blogs compared to conservative ones (panel (b)). Additionally, we report the truthful/strategic public opinion as a function of the logarithm of the eigenvector centrality $\bm{\pi}_{i}$ (cf. panels (c, d)) for each node, as well as the absolute change $|\bm{z}_{i}^{\prime}-\bm{z}_{i}|$ (cf. panel (e)). We fit a regression model, and we detect significant effects (${}^{***}:P<0.001,\;^{**}:P<0.01,\;^{*}:P<0.05$; effects denoted by $\theta$) of the logarithm of the centrality to the truthful equilibrium $\bm{z}$, the strategic equilibrium $\bm{z}^{\prime}$, and the change $|\bm{z}^{\prime}-\bm{z}|$, revealing the structure of a power law. Finally, we observe that relative changes are more dispersed along liberal sources compared to conservative sources (cf. panel (f)).
Optimal Deviation for One Agent and All Agents
Assuming that we have one strategic agent, what is the change in their opinion? We can show that the new opinion is a scalar multiple of the initial opinion plus a bias term, where neither the scalar multiple nor the bias term can be zero.
Corollary 2.6 (Deviation for One Agent).
Let $S=\{i\}$. Then, $\bm{s}_{i}^{\prime}=\theta_{i}\bm{s}_{i}+\beta_{i}$ where
$\theta_{i}$ $=\frac{\alpha_{i}B_{ii}}{(1-\alpha_{i})\sum_{i\sim j}w_{ij}(B_{ii}-B_{ij})^{2}+\alpha_{i}B_{ii}^{2}}>0,$
$\beta_{i}$ $=-\frac{\alpha_{i}\sum_{j\neq i}B_{ij}s_{j}}{(1-\alpha_{i})\sum_{i\sim j}w_{ij}(B_{ii}-B_{ij})^{2}+\alpha_{i}B_{ii}^{2}}.$
Similarly, we can relate the maximum deviation of $\bm{s}^{\prime}$ from $\bm{s}$ in the other extreme case, i.e., when all agents are deviating ($S=[n]$).
Corollary 2.7.
When all agents are deviating ($S=[n]$), and $\alpha_{i}=\alpha$, then $\bm{s}^{\prime}$ satisfies:
$\frac{\|\bm{s}^{\prime}\|_{2}}{\|\bm{s}\|_{2}}\leq\frac{\lambda_{n}+\tilde{\alpha}}{\tilde{\alpha}}.$
Proof of Corollary 2.7.
When all agents are deviating, it is straightforward to show that $\tilde{T}=\tilde{\alpha}B$ with minimum eigenvalue $\tilde{\alpha}^{2}/(\lambda_{n}+\tilde{\alpha})>0$. Thus $\tilde{T}$ is invertible, and therefore $\bm{s}^{\prime}=\frac{1}{\tilde{\alpha}}B^{-1}\widetilde{\textrm{diag}(B)}\bm{s}$, where $\widetilde{\textrm{diag}(B)}$ is a diagonal matrix with entries $B_{ii}$. Then
$\|\bm{s}^{\prime}\|_{2}$ $\leq\frac{1}{\tilde{\alpha}}\|B^{-1}\|_{2}\|\widetilde{\textrm{diag}(B)}\|_{2}\|\bm{s}\|_{2}$
$=\left(\max_{i}B_{ii}\right)\left(\max_{i}\frac{\lambda_{i}+\tilde{\alpha}}{\tilde{\alpha}}\right)\|\bm{s}\|_{2}$
$\leq\frac{\tilde{\alpha}}{\lambda_{1}+\tilde{\alpha}}\frac{\lambda_{n}+\tilde{\alpha}}{\alpha}\|\bm{s}\|_{2}$
$=\frac{\lambda_{n}+\tilde{\alpha}}{\tilde{\alpha}}\|\bm{s}\|_{2}.$
The proof of Corollary 2.7 shows that the adjusted susceptibility ($\tilde{\alpha}$) and the maximum eigenvalue of the Laplacian ($\lambda_{n}$) are responsible for changes in the norm of $\bm{s}^{\prime}$. From classic spectral graph theory, we know that $\lambda_{n}=\Theta(d_{\textup{max}})$ where $d_{\textup{max}}$ is the maximum degree of the graph; therefore, graphs with a lower maximum degree experience smaller distortions. Also, regarding the susceptibility to persuasion, the distortion becomes $1+o(1)$ as long as $\tilde{\alpha}=\omega(d_{\textup{max}})$.
Equilibria for real-world datasets. Next, we discuss the results of experiments simulating the strategically manipulated equilibria for our real-world datasets.
@@
Regarding real-world data, Figure 2 shows the relationship between the truthful opinions ($\bm{s}$ and $\bm{z}$) and the strategic ones ($\bm{s}^{\prime}$ and $\bm{z}^{\prime}$) for the datasets, along with the corresponding correlation coefficient $R^{2}$, assuming that $S$ consists of the top-50% nodes in terms of their eigenvector centrality, for susceptibility parameters set to $\alpha_{i}=0.5$ (equal self-persuasion and persuasion due to others) and $\alpha_{i}=0.25$ (higher persuasion due to others).
@@
Figure 3 analyzes the opinions of the strategic set $S$ on the Polblogs dataset. Specifically, we find that larger changes in sentiment happen across liberal outlets compared to conservative ones. Additionally, the changes in the truthful/strategic opinions are related to the eigenvector centrality $\bm{\pi}_{i}$ as a power law, i.e., $\bm{z}_{i}^{\prime}\propto\bm{\pi}_{i}^{\theta}$ ($P<0.001$; linear regression between the log centralities $\log\bm{\pi}_{i}$ and $\bm{z}_{i}^{\prime}$). The same finding holds for $|\bm{z}_{i}^{\prime}-\bm{z}_{i}|$ and $\bm{z}_{i}$.
At this point, one may wonder whether the eigenvector centrality really influences the strategic opinions $\bm{z}_{i}^{\prime}$ for $i\in S$. Our answer is negative. We repeat the same experiment with the Twitter and Reddit datasets, where we find no effects ($P>0.1$; linear regression between the log centralities $\log\bm{\pi}_{i}$ and $\bm{z}_{i}^{\prime}$). Due to space limitations, the corresponding figures are deferred to Appendix C.
@@
Figure 4 shows how the polarization, disagreement, and cost change as a function of the susceptibility parameter $\alpha_{i}$. Except for $\alpha_{i}\approx 0.3$, the polarization ratio, disagreement ratio, and the price of misreporting experience a downward trend as $\alpha_{i}$ increases. This indicates that as as users prioritize their own opinions more than their neighbors, they are less susceptible to strategic manipulation.
@@
Next, we study the effect of the number of deviators, which corresponds to $|S|$, on the changes in polarization, disagreement, and the total cost (through the price of misreporting). Figure 5 shows how the polarization and disagreement when $S$ consists of the top-1-10% most central agents with respect to eigenvector centrality. We show that even if only 1% of agents are strategic, this can impact consensus by several orders of magnitude.
@@ 3. Price of Misreporting
Theorem 3.1 shows that the PoM is small when the spectral radius of the Laplacian is small, and when agents are somewhat susceptible to their neigbhors ($\alpha\not\to 0$). Note that the spectral radius can be replaced by a degree bound: if $d_{\textup{max}}$ is the maximum degree of the graph, then $\lambda_{n}\leq 2d_{\textup{max}}$. So the PoM is small if the maximum degree is small.
@@
$\mathsf{PoM}\leq\frac{(\lambda_{n}+4\tilde{\alpha})(\lambda_{n}+\tilde{\alpha})^{2}}{\tilde{\alpha}^{5}}=O\left(\max\left\{\frac{\lambda_{n}}{\tilde{\alpha}^{5}},\frac{1}{\tilde{\alpha}^{2}}\right\}\right).$
Proof of Theorem 3.1..
(11) $Q=BLB+\tilde{\alpha}(I-2B+B^{2})$
Since $Q\succeq 0$, it has eigendecomposition $Q=U\Lambda_{Q}U^{T}$. Moreover, $U$ is precisely the matrix of eigenvectors for the Laplacian. The eigenvalues of $Q$ can be shown to be $\tilde{\alpha}^{2}/(\lambda_{i}+\tilde{\alpha})$. Therefore, $C(z)=\bm{s}^{T}Q\bm{s}\geq\frac{\tilde{\alpha}^{2}}{\lambda_{n}+\tilde{\alpha}}\|\bm{s}\|_{2}^{2}$.
Next, let $\textrm{diag}(B)$ be the diagonal matrix with entries $B_{ii}$ and $\widetilde{\textrm{diag}(B)}$ be as in Corollary 2.7. In the proof of Corollary 2.7, we show that $\bm{s}^{\prime}=(1/\tilde{\alpha})B^{-1}\widetilde{\textrm{diag}(B)}\bm{s}$ and $\bm{z}^{\prime}=(1/\tilde{\alpha})\widetilde{\textrm{diag}(B)}\bm{s}$, which similarly implies (after algebraic operations) that $C(\bm{z}^{\prime})/(1-\alpha)=\bm{s}^{T}Q^{\prime}\bm{s}$ where:
$Q^{\prime}:=\frac{1}{\tilde{\alpha}^{2}}\widetilde{\textrm{diag}(B)}L\widetilde{\textrm{diag}(B)}+\frac{1}{\tilde{\alpha}}B^{-1}\left(\widetilde{\textrm{diag}(B)}\right)^{2}B^{-1}-2\frac{1}{\tilde{\alpha}}B^{-1}\left(\widetilde{\textrm{diag}(B)}\right)^{2}+\frac{1}{\tilde{\alpha}}\left(\widetilde{\textrm{diag}(B)}\right)^{2}$
(12) $\|\widetilde{\textrm{diag}(B)}\|_{2}$ $=\max_{i}B_{ii}\leq\|B\|_{2}=1,$
By the triangle inequality, the Cauchy-Schwarz inequality, and Equations 12 and 13, we have that:
$\|Q^{\prime}\|_{2}$ $\leq\frac{1}{\tilde{\alpha}^{2}}\|L\|_{2}\left(\|\widetilde{\textrm{diag}(B)}\|_{2}\right)^{2}+\frac{1}{\tilde{\alpha}}\left(\|\widetilde{\textrm{diag}(B)}\|_{2}\right)^{2}\|B^{-1}\|_{2}^{2}+\frac{2}{\tilde{\alpha}}\|B^{-1}\|_{2}+\frac{1}{\tilde{\alpha}}\left(\|\widetilde{\textrm{diag}(B)}\|_{2}\right)^{2}$
Finally, we can simplify:
Corollary 3.2 (Price of Misreporting for Heterogeneous Susceptibility).
Proof.
Note that $C(\bm{z}^{\prime})\leq(1-\alpha_{\min})(\bm{z}^{\prime})^{T}L\bm{z}^{\prime}+\alpha_{\max}\|\bm{z}^{\prime}-\bm{s}\|_{2}^{2}=\overline{C}(\bm{z}^{\prime})$, and $C(\bm{z})\geq(1-\alpha_{\max})\bm{z}^{T}L\bm{z}+\alpha_{\min}\|\bm{z}-\bm{s}\|_{2}^{2}=\underline{C}(\bm{z})$ where $\alpha_{\min}=\min_{i\in[n]}\alpha_{i}$, and $\alpha_{\max}=\max_{i\in[n]}\alpha_{i}$. Then, the same analysis of Theorem 3.1 can be applied, since $\overline{C}(\bm{z}^{\prime})/\underline{C}(\bm{z})\geq C(\bm{z}^{\prime})/C(\bm{z})$. ∎
Finally, we discuss how one may generalize Theorem 3.1 to the case where some agents are honest.
Towards fine-grained $\mathsf{PoM}$ guarantees.
Figure 4 shows that the PoM is not monotonic in $\lvert S\rvert$. As the number of strategic agents grows, the PoM can fall or grow, depending on the choice of $S$, network parameters, and so on. Therefore, we would like to give a version of Theorem 3.1 for any set of strategic agents $S\subset[n]$, not just the case of $S=[n]$. However, proving such a bound would require analyzing $S\times S$ principal submatrices of $B,L$ to obtain characterizations of the cost at the corrupted equilibrium $\bm{z}^{\prime}$. In particular, we would require a restricted invertibility estimate to prove the analogue of Eq. (13). To our knowledge, the best such estimates (Marcus et al., 2022) are too lossy when $n-\lvert S\rvert$ is large. We leave this question to future work.
@@ 4. Learning from Network Outcomes
4.1. Detecting Manipulation with a Hypothesis Test
In many real-world networks, the distribution of truthful opinions follows a Gaussian distribution (Figure 6). Given estimates $(\widehat{L},\widehat{A})$ for the graph Laplacian and susceptibility matrix, the platform can observe the corrupted equilibrium $\bm{z}^{\prime}$ and solve for the strategic opinions $\bm{s}^{\prime}$ via:
@@ 4.1. Detecting Manipulation with a Hypothesis Test
We propose that the platform perform a one-sample $t$-test with the entries $\bm{s}^{\prime}$, with a population mean $\mu_{0}\in\mathbb{R}$ based on e.g. historical data. Under the null hypothesis in which no manipulation has occurred, $\bm{s}^{\prime}=\bm{s}$, so the test should fail to reject the null hypothesis. However, when agents $S\subset[n]$ deviate, then $\bm{s}_{i}^{\prime}\neq\bm{s}_{i}$ for $i\in S$, so the test should reject the null for large enough deviations. The test is simple, and described in Algorithm 1. Figure 6 shows the results of the test for varying choices of $S$. We see that at significance level $0.05$, the test has low Type I error, as it will return “No Manipulation” when $S=\emptyset$, and low Type II error as it will return “Manipulation” when $S\neq\emptyset$.
@@
For the Political Blogs dataset, intrinsic opinions belong to {±$\{\pm 1\}$, so the null hypothesis should be a biased Rademacher distribution. In this case, one should use a $\chi^{2}$-test, as in  Agarwal et al. (2020); Chen and Tsourakakis (2022).
Figure 6. The true opinions for Twitter (left) and Reddit (middle) both follow a normal distribution. When simulating strategic manipulation with random choices of $S$ (right), the detection test (Algorithm 1) has no Type I or Type II error at significance level $0.05$. The tester uses $\widehat{L}=L,\widehat{A}=A=\frac{1}{2}I$, and $\mu_{0}$ equal to the mean of the true intrinsic opinions. Shaded regions are $95\%$ confidence intervals for $p$-values of the test across $5$ independent runs.
4.2. Learning the Strategic Actors with Robust Regression
The algorithm described in the previous section (Algorithm 1) can be used to detect whether there exists manipulation in the network. However, the set $S$ is unknown, and therefore the platform cannot target the deviators to perform interventions to mitigate strategic behavior.
It is, therefore, essential for the platform to be able to identify the set of deviators $S$, in case the platform needs to take regulatory actions. While at first, it may seem that finding the set of deviators $S$ is a hard task, it turns out that under mild assumptions on the intrinsic opinion formation process, we can learn the set of deviators $S$ from observing the strategically corrupted equilibrium $\bm{z}^{\prime}$ via Algorithm 2 in polynomial time, described in Algorithm 2. Our algorithm is based on robust regression leveraging the Torrent algorithm developed by Bhatia et al. (2015) and requires access to a node embedding matrix $X\in\mathbb{R}^{n\times d}$, and the size $|S|$ of the set of deviators.
@@ 4.2. Learning the Strategic Actors with Robust Regression
The key idea of Algorithm 2 is that if the size of the strategic set $S$ is sufficiently small, in general, $|S|\leq Cn$ for some small constant $C$, then we can view the misreported intrinsic opinions $\bm{s}^{\prime}$ as a perturbation of the truthful opinion vector $\bm{s}$, and then use a robust regression algorithm to estimate $\bm{s}$. We assume that the embedding matrix $X\in\mathbb{R}^{n\times d}$ determines intrinsic beliefs: for example, demography, geographic location, etc. Node-level features can be learned by a variety of methods, such as spectral embeddings on the graph Laplacian or graph neural networks. Previous works have used the framework of combining a robust estimator with model-specific information to learn from “strategic sources” of data, such as in bandits (Kapoor et al., 2019), controls (Russo, 2023), and network formation games (Jalan and Chakrabarti, 2024).
@@
We give our guarantee for Algorithm 2.
@@
Let $X$ be as in Algorithm 2, and suppose that $X\bm{v}=\bm{s}$ for some $\bm{v}\in\mathbb{R}^{d}$, and that $X$ satisfies the SSC condition at level $1-\gamma$ with constant $\xi_{1-\gamma}$, and SSS condition at level $\gamma$ with constant $\Xi_{\gamma}$ (Definition 4.1). Then, there exist absolute constants $C,C^{\prime}>0$ such that if $\lvert S\rvert\leq Cn$ and $14\frac{\sqrt{\Xi_{\gamma}}}{\sqrt{\xi_{1-\gamma}}}<$ 1, Algorithm 2 returns $\hat{\bm{s}}$ such that:
$\|\hat{\bm{s}}-\bm{s}\|_{2}\leq\|X\|_{2}n^{-\omega(1)},$
using $T=C^{\prime}(\log n)^{2}$ iterations of Torrent for the Robust Regression step. Moreover, if for all $j\in S$ we have $\lvert\bm{s}_{j}-\bm{s}_{j}^{\prime}\rvert\gg\|X\|_{2}n^{-\omega(1)}$, then $\hat{S}=S$.
@@
We are ready to prove Proposition 4.2.
Proof of Proposition 4.2.
Let $\widehat{\bm{s}^{\prime}}$ be as in Algorithm 2, and $\bm{y}=\widehat{\bm{s}^{\prime}}$. Notice $\bm{y}=\bm{s}+(\bm{s}^{\prime}-\bm{s})+(\widehat{\bm{s}^{\prime}}-\bm{s}^{\prime})$. Let $\bm{w}:=(\bm{s}^{\prime}-\bm{s})$ be the corruption vector due to strategic negotiations and $\bm{r}=(\widehat{\bm{s}^{\prime}}-\bm{s}^{\prime})$ be the residual vector due to least-squares regression. We claim that $\bm{r}=\bm{0}$, because $A^{-1}$ is full rank and $((I-A)L+A)$ is full rank, so $\widehat{\bm{s}^{\prime}}=A^{-1}((I-A)L+A)\bm{z}^{\prime}=\bm{s}^{\prime}$.
Therefore, letting $T=C^{\prime}(\log n)^{2}$ for large enough constant $C^{\prime}>0$, we see that $\|\hat{\bm{v}}-\bm{v}\|_{2}\leq O(n^{-\omega(1)})$. Hence $\|\hat{\bm{s}}-\bm{s}\|_{2}=\|X\hat{\bm{v}}-X\bm{v}\|_{2}\leq\|X\|_{2}\cdot n^{-\omega(1)}$. Now, let $\bm{u}=\hat{\bm{s}}-\bm{s}^{\prime}$. If $i\in[n]\setminus S$, then $\lvert\bm{u}_{i}\rvert\leq\|X\|_{2}\cdot n^{-\omega(1)}$. On the other hand for $j\in S$, $\lvert\bm{u}_{j}\rvert\geq\lvert\bm{s}_{j}-\bm{s}_{j}^{\prime}\rvert-\|X\|_{2}n^{-\omega(1)}$. Therefore the top-$|S|$ entries of $\bm{u}$ recover $S$. ∎
@@
If $G$ has two communities with $n_{1}$ and $n_{2}$ nodes respectively such that $n_{1}\geq n_{2}\geq 1$, and $X\in\{0,1\}^{n\times 2}$ is an embedding vector where each row $\bm{x}_{i}$ corresponds to a one-hot vector for the community of node $i$, then Algorithm 2 can recover $S$ perfectly as long as $|S|<\frac{n_{1}}{17}$.
For instance, when Proposition 4.4 is applied to the Polblogs dataset, it shows that $S$ can be fully recovered as long as $|S|\leq 9$. We can obtain a slightly worse bound and extend the result to a blockmodel graph with $K$ communities (proof in the Appendix).
@@
If $G$ has $K\geq 2$ communities with sizes $n_{1}\geq n_{2}\geq\dots\geq n_{K}$ with $\left(\frac{16K}{16K+1}\right)\frac{n}{K}<n_{K}\leq\frac{n}{K}$ and $X\in\{0,1\}^{n\times K}$ is an embedding vector where $\bm{x}_{i}$ corresponds to an one-hot encoding of the community membership, then Algorithm 2 can recover $S$ perfectly as long as $|S|<\frac{1}{16K+1}n$.
In detail, Proposition 4.5 states that as long as the smallest community of the graph has size $\Theta(n/K)$ then the recovery of a set of size $|S|=O(n/K)$ is possible. If $\lvert S\rvert\gg n/K$ and $S$ contains all members of the smallest community, then robust regression can fail.
@@
We show that in real-world datasets, Algorithm 2 can identify the set of deviators with high accuracy (cf. Figure 7). Specifically, in the real-world datasets we take $S$ to be a randomly sampled set of size $\lceil pn\rceil$ for $p\in\{0.05,0.1,0.15,0.2\}$ and the embeddings to be 128-dimenaional Node2Vec embeddings for Twitter and Reddit and community membership embeddings for the Pollblogs dataset. Algorithm 2 achieves low recovery error as well as high balanced accuracy score.

### v2

@@ 2. Strategic Opinion Formation
Figure 3. Strategic misreports on the Polblogs dataset, where $S$ consists of the top 50% of agents by eigenvector centrality. Nodes are labeled as liberal ($\bm{s}_{i}=-1$) or conservative ($\bm{s}_{i}=+1$). A node is said to change belief if $\bm{z}_{i}$ and $\bm{z}_{i}^{\prime}$ have opposite signs. In scatterplots (a), (c), (d), and (e), point shape indicates whether a belief change occurs, while color denotes the intrinsic opinion. We observe more frequent belief changes among liberal blogs than conservative ones (panel (b)). Panels (c) and (d) report truthful and strategic public opinions as functions of $\log\bm{\pi}_{i}$, while panel (e) shows the absolute change $|\bm{z}_{i}^{\prime}-\bm{z}_{i}|$. Regression analysis reveals significant effects of $\log\bm{\pi}_{i}$ ($^{***}\!:\,P<0.001$, $^{**}\!:\,P<0.01$, $^{*}\!:\,P<0.05$; coefficients $\theta$) on $\bm{z}$, $\bm{z}^{\prime}$, and $|\bm{z}^{\prime}-\bm{z}|$, consistent with power-law structure. Finally, relative belief changes are more dispersed among liberal sources than conservative ones (panel (f)).
@@
Regarding real-world data, Figure˜2 shows the relationship between the truthful opinions ($\bm{s}$ and $\bm{z}$) and the strategic ones ($\bm{s}^{\prime}$ and $\bm{z}^{\prime}$) for the datasets, along with the corresponding correlation coefficient $R^{2}$, assuming that $S$ consists of the top-50% nodes in terms of their eigenvector centrality, for susceptibility parameters set to $\alpha_{i}=0.5$ (equal self-persuasion and persuasion due to others) and $\alpha_{i}=0.25$ (higher persuasion due to others).
@@
Figure˜3 analyzes the opinions of the strategic set $S$ on the Polblogs dataset. Specifically, we find that larger changes in sentiment happen across liberal outlets compared to conservative ones. Additionally, the changes in the truthful/strategic opinions are related to the eigenvector centrality $\bm{\pi}_{i}$ as a power law, i.e., $\bm{z}_{i}^{\prime}\propto\bm{\pi}_{i}^{\theta}$ ($P<0.001$; linear regression between the log centralities $\log\bm{\pi}_{i}$ and $\bm{z}_{i}^{\prime}$). The same finding holds for $|\bm{z}_{i}^{\prime}-\bm{z}_{i}|$ and $\bm{z}_{i}$.
At this point, one may wonder whether the eigenvector centrality really influences the strategic opinions $\bm{z}_{i}^{\prime}$ for $i\in S$. Our answer is negative. We repeat the same experiment with the Twitter and Reddit datasets, where we find no effects ($P>0.1$; linear regression between the log centralities $\log\bm{\pi}_{i}$ and $\bm{z}_{i}^{\prime}$). Due to space limitations, the corresponding figures are deferred to Appendix˜B.
@@
Figure˜4 shows how the polarization, disagreement, and cost change as a function of the susceptibility parameter $\alpha_{i}$. Except for $\alpha_{i}\approx 0.3$, the polarization ratio, disagreement ratio, and the price of misreporting experience a downward trend as $\alpha_{i}$ increases. This indicates that as as users prioritize their own opinions more than their neighbors, they are less susceptible to strategic manipulation.
@@
Next, we study the effect of the number of deviators, which corresponds to $|S|$, on the changes in polarization, disagreement, and the total cost (through the price of misreporting). Figure˜5 shows how the polarization and disagreement when $S$ consists of the top-1-10% most central agents with respect to eigenvector centrality. We show that even if only 1% of agents are strategic, this can impact consensus by several orders of magnitude.
@@ 3. Price of Misreporting
Theorem 3.1 shows that the PoM is small when the spectral radius of the Laplacian is small, and when agents are somewhat susceptible to their neighbors ($\alpha\not\to 0$). Note that the spectral radius can be replaced by a degree bound: if $d_{\textup{max}}$ is the maximum degree of the graph, then $\lambda_{n}\leq 2d_{\textup{max}}$. So the PoM is small if the maximum degree is small.
@@
$\mathsf{PoM}\leq\frac{(\lambda_{n}+4\tilde{\alpha})(\lambda_{n}+\tilde{\alpha})^{2}}{\tilde{\alpha}^{5}}=O\left(\max\left\{\frac{\lambda_{n}^{3}}{\tilde{\alpha}^{5}},\frac{1}{\tilde{\alpha}^{2}}\right\}\right).$
Next, we discuss how one may generalize Theorem 3.1 to the case where some agents are honest: Figure 4 shows that the PoM is not monotonic in $\lvert S\rvert$. As the number of strategic agents grows, the PoM can fall or grow, depending on the choice of $S$, network parameters, and so on. Therefore, we would like to give a version of Theorem 3.1 for any set of strategic agents $S\subset[n]$, not just the case of $S=[n]$. However, proving such a bound would require analyzing $S\times S$ principal submatrices of $B,L$ to obtain characterizations of the cost at the corrupted equilibrium $\bm{z}^{\prime}$. In particular, we would require a restricted invertibility estimate to prove the analogue of Eq. (13). To our knowledge, the best such estimates (Marcus et al., 2022) are too lossy when $n-\lvert S\rvert$ is large. We leave this question to future work.
@@ 4. Learning from Network Outcomes
Detecting Manipulation with a Hypothesis Test
[…]
@@
We propose that the platform perform a one-sample $t$-test with the entries $\bm{s}^{\prime}$, with a population mean $\mu_{0}\in\mathbb{R}$ based on e.g. historical data. Under the null hypothesis in which no manipulation has occurred, $\bm{s}^{\prime}=\bm{s}$, so the test should fail to reject the null hypothesis. However, when agents $S\subset[n]$ deviate, then $\bm{s}_{i}^{\prime}\neq\bm{s}_{i}$ for $i\in S$, so the test should reject the null for large enough deviations. The test is simple, and described in Algorithm˜1. Figure 6 shows the results of the test for varying choices of $S$. We see that at significance level $0.05$, the test has low Type I error, as it will return “No Manipulation” when $S=\emptyset$, and low Type II error as it will return “Manipulation” when $S\neq\emptyset$.
@@
For the Political Blogs dataset, intrinsic opinions belong to $\{\pm 1\}$, so the null hypothesis should be a biased Rademacher distribution. In this case, one should use a $\chi^{2}$-test, as in  Agarwal et al. (2020); Chen and Tsourakakis (2022).
Figure 6. The true opinions for Twitter (left) and Reddit (middle) both follow a normal distribution. When simulating strategic manipulation with random choices of $S$ (right), the detection test (Algorithm 1) has no Type I or Type II error at significance level $0.05$. The tester uses $\widehat{L}=L,\widehat{A}=A=\frac{1}{2}I$, and $\mu_{0}$ equal to the mean of the true intrinsic opinions. Shaded regions are $95\%$ confidence intervals for $p$-values of the test across 55 independent runs.
Learning the Strategic Actors with Robust Regression
The algorithm described in the previous section (Algorithm˜1) can be used to detect whether there exists manipulation in the network. However, the set $S$ is unknown, and therefore the platform cannot target the deviators to perform interventions to mitigate strategic behavior.
It is, therefore, essential for the platform to be able to identify the set of deviators $S$, in case the platform needs to take regulatory actions. While at first, it may seem that finding the set of deviators $S$ is a hard task, it turns out that under mild assumptions on the intrinsic opinion formation process, we can learn the set of deviators $S$ from observing the strategically corrupted equilibrium $\bm{z}^{\prime}$ via Algorithm˜2 in polynomial time, described in Algorithm˜2. Our algorithm is based on robust regression leveraging the Torrent algorithm developed by Bhatia et al. (2015) and requires access to a node embedding matrix $X\in\mathbb{R}^{n\times d}$, and the size $|S|$ of the set of deviators.
@@
The key idea of Algorithm˜2 is that if the size of the strategic set $S$ is sufficiently small, in general, $|S|\leq Cn$ for some small constant $C$, then we can view the misreported intrinsic opinions $\bm{s}^{\prime}$ as a perturbation of the truthful opinion vector $\bm{s}$, and then use a robust regression algorithm to estimate $\bm{s}$. We assume that the embedding matrix $X\in\mathbb{R}^{n\times d}$ determines intrinsic beliefs: for example, demography, geographic location, etc. Node-level features can be learned by a variety of methods, such as spectral embeddings on the graph Laplacian or graph neural networks. Previous works have used the framework of combining a robust estimator with model-specific information to learn from “strategic sources” of data, such as in bandits (Kapoor et al., 2019), controls (Russo, 2023), and network formation games (Jalan and Chakrabarti, 2024).
@@
We give our guarantee for Algorithm˜2.
@@
Let $X$ be as in Algorithm 2, and suppose that $X\bm{v}=\bm{s}$ for some $\bm{v}\in\mathbb{R}^{d}$, and that $X$ satisfies the SSC condition at level $1-\gamma$ with constant $\xi_{1-\gamma}$, and SSS condition at level $\gamma$ with constant $\Xi_{\gamma}$ (Definition 4.1). Then, there exist absolute constants $C,C^{\prime}>0$ such that if $\lvert S\rvert\leq Cn$ and $4\frac{\sqrt{\Xi_{\gamma}}}{\sqrt{\xi_{1-\gamma}}}<1$, Algorithm˜2 returns $\hat{\bm{s}}$ such that $\|\hat{\bm{s}}-\bm{s}\|_{2}\leq\|X\|_{2}n^{-\omega(1)}$, using $T=C^{\prime}(\log n)^{2}$ iterations of Torrent for the Robust Regression step. Moreover, if for all $j\in S$ we have $\lvert\bm{s}_{j}-\bm{s}_{j}^{\prime}\rvert\gg\|X\|_{2}n^{-\omega(1)}$, then $\hat{S}=S$.
@@
If $G$ has two communities with $n_{1}$ and $n_{2}$ nodes respectively such that $n_{1}\geq n_{2}\geq 1$, and $X\in\{0,1\}^{n\times 2}$ is an embedding vector where each row $\bm{x}_{i}$ corresponds to a one-hot vector for the community of node $i$, then Algorithm˜2 can recover $S$ perfectly as long as $|S|<\frac{n_{1}}{17}$.
For instance, when Proposition˜4.4 is applied to the Polblogs dataset, it shows that $S$ can be fully recovered as long as $|S|\leq 9$. We can obtain a slightly worse bound and extend the result to a blockmodel graph with $K$ communities (proof in the Appendix).
@@
If $G$ has $K\geq 2$ communities with sizes $n_{1}\geq n_{2}\geq\dots\geq n_{K}$ with $\left(\frac{16K}{16K+1}\right)\frac{n}{K}<n_{K}\leq\frac{n}{K}$ and $X\in\{0,1\}^{n\times K}$ is an embedding vector where $\bm{x}_{i}$ corresponds to an one-hot encoding of the community membership, then Algorithm˜2 can recover $S$ perfectly as long as $|S|<\frac{1}{16K+1}n$.
@@
In detail, Proposition˜4.5 states that as long as the smallest community of the graph has size $\Theta(n/K)$ then the recovery of a set of size $|S|=O(n/K)$ is possible. If $\lvert S\rvert\gg n/K$ and $S$ contains all members of the smallest community, then robust regression can fail. We show that in real-world datasets, Algorithm˜2 can identify the set of deviators with high accuracy (cf. Figure˜7). Specifically, in the real-world datasets we take $S$ to be a randomly sampled set of size $\lceil pn\rceil$ for $p\in\{0.05,0.1,0.15,0.2\}$ and the embeddings to be 128-dimenaional Node2Vec embeddings for Twitter and Reddit and community membership embeddings for the Pollblogs dataset. Algorithm˜2 achieves low recovery error as well as high balanced accuracy score.